OpenStudy (anonymous):

Calculate the wavelength of a photon of energy 5.518 × 10^-26 joules. (Planck's constant is 6.626 × 10^-34 joule seconds, the speed of light is 2.998 × 10(m/s)

OpenStudy (anonymous):

E = h * ni( frequency) c=ni * lambda(wavelenght)

OpenStudy (anonymous):

and h is Plancks constant

OpenStudy (anonymous):

lambda (wavelength) = (h* c)/ E = 3,5999 ~ 3,6 m